OpenStudy (anonymous):

Ravi and his wife are each starting a saving plan. Ravi will initially set aside $50 and then add $40.75 every week to the savings. The amount A (in dollars) saved this way is given by the function A = 40.75N + 50, where N is the number of weeks he has been saving. His wife will not set an initial amount aside but will add $75.65 to the savings every week. The amount B (in dollars) saved using this plan is given by the function B = 75.65N. Let T be total amount (in dollars) saved using both plans combined. Write an equation relating T to N. Simplify your answer as much as possible.

OpenStudy (er.mohd.amir):

simply add both amounts get saving by both

OpenStudy (anonymous):

"simply add both amounts get saving by both" Are you drunk or am I? lmao

OpenStudy (anonymous):

I really need help. what exactly do I add?

OpenStudy (er.mohd.amir):

40.75N+50 and 75.65N

OpenStudy (anonymous):

So my answer would be 166.4?

OpenStudy (er.mohd.amir):

add same type that is N with N and other withh other

OpenStudy (anonymous):

I am not understanding you.

OpenStudy (anonymous):

@legendarysadist what do you think?

OpenStudy (anonymous):

ARe you saying to add, 40.75 + 40.75 and 75.65 + 75.65?

OpenStudy (er.mohd.amir):

40.75N+75.65N=(40.75+75.65)N=?

OpenStudy (anonymous):

Well the amount they save is detailed in the plans A and B. So if you want to find the total amount saved as a function of T, it would be that T=A+B. Their two plans of saving money added together will be the total amount saved. But it says that it wants T to be related to N ,not A or B. So you have to add the equivalents of A and B, which are "40.75N+50" and "75.65N".

OpenStudy (anonymous):

OK so 40.75 + 50 + 75.65

OpenStudy (er.mohd.amir):

No

OpenStudy (anonymous):

You have to keep the N's. You don't add the 50 with the others, they are of the separate value, N. It would be more like (40.75N+75.65N)+50

OpenStudy (anonymous):

Like terms are added together.

OpenStudy (anonymous):

116.4n + 50 would be the answer?

OpenStudy (anonymous):

@er.mohd.amir @legendarysadist

OpenStudy (anonymous):

Well yes, but this is represented as a function of T, so it should be written out as \[\large \sf T=116.4N+50\]

OpenStudy (er.mohd.amir):

yes use N instant of n

OpenStudy (anonymous):

OK thank you so much! You have no idea how much you have helped!

OpenStudy (anonymous):

Ye no problem (:

OpenStudy (anonymous):

I have one more. Would you be able to help?